OpenStudy (anonymous):

the stopping distance d of a car after the brakes are applied varies directly as the square of speed r. If a car traveling 40 mph can stop in 90 ft how many feet will it take the same car to stop when it is traveling 60 mph

OpenStudy (anonymous):

202.5 feet

OpenStudy (dumbcow):

d = kr^2 90 = k(40)^2 k = 90/1600 = 9/160 d = (9/160)(60^2) d = 202.5